Please can anyone help me I am lost.
17. On three consecutive passes, a football team gains 9 yards, loses 16 yards, and gains 42 yards. What number represents the total net yardage?
The total net yardage is _______yds.
18. Use the distributive property to help you solve the equation.
3(w-3)=-9
w=______
19. Find the domain of the function.
g(x)= 6/8-9x
a. {x|x is a real number and x >= 8/9} that is greater than or equal too
b. {x|x is a real number and x =/ 0} the = should have the / thru it.
c. {x|x is a real number and x=/ 8/9}
d. {x|x is a real number and x =/ 6}
21. Evaluate
r+s/ 8 for r= 40 and s=16
r+s/8= _______
22. Find the domain of the function.
p(x)= x to the 2 power - 2x +7
A. {x|x is a real number and x =/0}
b. {x|x is a real number and x =/7}
c. {x|x is a real number}
d. {x|x is a real number and x>0}
24. Solve using the multiplication principle. Dont forget to perform a check.
12x= -84
The solution is x= _______?



I thank you in advance for you knowledge and generousity in helping people like me who are mathmatically challenged.
Deborah

Answer by JBarnum(2146)

17. On three consecutive passes, a football team gains 9 yards, loses 16 yards, and gains 42 yards. What number represents the total net yardage?
The total net yardage is _______yds.
9-16+42 => -5+42 => 37yards
18. Use the distributive property to help you solve the equation.
3(w-3)=-9
w=______
(3)w -(3)3=-9 => 3w-9=-9 => 3w=0 => w=0,
check: 3(0-3)=-9 => 3(-3)=-9 => -9=-9
21. Evaluate
r+s/ 8 for r= 40 and s=16
r+s/8= _______
(40+16)/8 => 56/8=7
24. Solve using the multiplication principle. Dont forget to perform a check.
12x= -84
The solution is x= _______?
-84/12=x => -7=x, 12(-7)=-84 => -84=-84